### v1

# Intercept Cancer: Cancer Pre-Screening with Large Scale Healthcare Foundation Models

Abstract
Cancer screening, leading to early detection, saves lives. Unfortunately, existing screening techniques require expensive and intrusive medical procedures, not globally available, resulting in too many lost would-be-saved lives. We present CATCH-FM, CATch Cancer early with Healthcare Foundation Models, a cancer pre-screening methodology that identifies high-risk patients for further screening solely based on their historical medical records. With millions of electronic healthcare records (EHR), we establish the scaling law of EHR foundation models pretrained on medical code sequences, pretrain compute-optimal foundation models of up to 2.4 billion parameters, and finetune them on clinician-curated cancer risk prediction cohorts. In our retrospective evaluation comprising of thirty thousand patients, CATCH-FM achieved strong efficacy (60% sensitivity) with low risk (99% specificity and Negative Predictive Value), outperforming feature-based tree models as well as general and medical large language models by large margins. Despite significant demographic, healthcare system, and EHR coding differences, CATCH-FM achieves state-of-the-art pancreatic cancer risk prediction on the EHRSHOT few-shot leaderboard, outperforming EHR foundation models pretrained using on-site patient data. Our analysis demonstrates the robustness of CATCH-FM in various patient distributions, the benefits of operating in the ICD code space, and its ability to capture non-trivial cancer risk factors. Our code will be open-sourced.
1 Introduction
Early cancer detection by cancer screening is one of the most effective ways to combat cancer [71]. Cancers detected at an early stage are treated with significantly improved patient outcomes [32; 46]. Recent medical advancements also significantly improved the curative rates for cancers detected at early stages [42; 61; 74; 85]. Routine screening with follow-up monitoring of high cancer risk patients is standard practice, enabling timely intervention and effective treatments [2; 31; 68].
Despite its benefits, cancer screening remains underutilized [105], especially in populations with limited healthcare resources [99] due to reliance on invasive, resource-intensive procedures like medical imaging [67; 96]. Screening is more common for cancers with clear risk factors, like breast and colorectal [72], while cancers without early symptoms, such as pancreatic cancer, often progress silently and are detected late, with survival durations under one year [5].
We present CATCH-FM: CATch Cancer early with Healthcare Foundation Models, a cancer pre-screening tool that identifies high-risk patients using only their medical history. CATCH-FM is pretrained on large-scale longitudinal EHR data and finetuned on clinician-curated cancer cohorts. It directly operates on precise medical codes (ICD), learning general medical patterns through next-code prediction and finetuned to capture cancer risk signals reflected in patient medical history [49; 50; 63]. Once trained, CATCH-FM can be deployed in EHR systems to predict cancer risk at low cost, supporting healthcare providers in deciding when and whom to screen.
To facilitate the study of EHR foundation models in cancer pre-screening, we build NHIRD-Cancer, a cancer risk prediction benchmark, by sampling more than three million patients from the Taiwanese National Health Insurance Research Database (NHIRD) [35], a government de-identified and research-accessible EHR database. We allocate three million patients for pretraining, consisting of billions of medical events spanning two decades, and the remainder for finetuning and evaluation. We focus on three cancers based on their critical needs for early detection: pancreatic, liver, and lung cancers [46; 48; 85], and curate a clinically reliable training and testing set by matching cancer patients with control-to-case group patient cohorts.
With billions of medical events available, we examine how compute budget (FLOPs), model size, and pretraining tokens affect performance. Our findings establish an EHR foundation model scaling law and confirm the benefit of large-scale pretraining on EHR data. Accordingly, we pretrain compute-optimal CATCH-FM models up to 2.4b parameters. When finetuned and evaluated for cancer risk prediction on NHIRD-Cancer, CATCH-FM consistently outperforms feature-based tree models and language models trained on the same data. Without loss of flexibility, it achieves strong high-risk prediction efficacy (60%+ sensitivity) with reliable classification of low-risk patients (99% specificity and Negative Predictive Value), often considered a minimum standard for screening techniques [7; 28]).
On the pancreatic cancer risk prediction task from Stanford Medicine EHRSHOT benchmark [97], CATCH-FM outperforms the prior state-of-the-art EHR foundation model, CLMBR [97], pretrained on millions of on-site patient records, while CATCH-FM is trained on data from drastically different populations, disease prevalence, and coding systems (ICD vs. SNOMED). Our analyses further demonstrate the robustness of CATCH-FM across different patient cohorts, preexisting conditions, and pre-screening configurations. Our interpretability analyses [27] reveal that CATCH-FM identified not only known cancer risk factors but also non-trivial markers discovered in recent medical research.
We view CATCH-FM as an effective, low-risk, and widely performable pre-screening tool that can assist healthcare providers make informed, effective, and efficient cancer screening decisions. To facilitate future research and development, our data curation and modeling code will be open-sourced under MIT license. The NHIRD-Cancer benchmark and trained model checkpoints will be released under the same license as NHIRD, enabling reproducibility and future research within the necessary constraints of privacy and regulations on patient data.
2 Related Work
Cancer screening significantly improves prognosis and patient outcomes [46; 48; 85]. Advancements in cancer treatment have made cancers more treatable and potentially curable, if they are detected in early stages [16; 36; 46; 55]. Recent advances in AI-powered medical imaging, like CT scans with multimodal models, have enhanced cancer screening accuracy, sometimes exceeding human-level sensitivity [7; 11; 94; 98]. The challenge is that medical imaging is resource-intensive, inaccessible for under-served populations, and not widely performable [23; 87; 90; 91].
Using electronic health records (EHR) to assess cancer risk is a promising path to improve cancer screening effectiveness and efficiency, i.e., by identifying patients with high cancer risk for healthcare professionals to make informed cancer screening decisions [49; 50]. Recent approaches have made attempts using feature-based machine learning models on large, task-specific datasets to detect cancer risks, and have shown the possibility of capturing cancer risk signals in EHR [62; 64].
Large language models pretrained on medical corpora are effective on text-oriented medical tasks [17] such as medical question answering [73], clinical document summarization [57], and radiology report generation [80]. Recent research has also pretrained foundation models on structured EHR data [15; 26; 103], and explored their utilities in various clinical tasks [70; 79; 102]. A challenge for EHR foundation models is that many EHR datasets only cover a slice of patients’ healthcare records [24; 41; 66], e.g., MIMIC-IV mainly focuses on emergency departments and ICU encounters [41]. Foundation models pretrained on EHR slices are more effective in corresponding prediction tasks such as emergency department triage [79] and ICU readmission risk [40].
Many tasks, like cancer risk prediction, require longitudinal EHR data to capture full patient history. Foresight uses two decades of data to pretrain Transformers for biomedical forecasting [47]. MOTOR trains a 143M-parameter model on time-to-event tasks using 2.7M patient records from 2014–2022, effectively predicting diagnosis time [75]. CLMBR, a 141M-parameter Transformer, is pretrained on 2.57M Stanford Medicine records over three decades [97], with 6.7K records released in the EHRSHOT benchmark for 15 few-shot prediction tasks [97].
Due to privacy and ethical constraints, large-scale longitudinal EHR data are hard to release publicly. Most EHR foundation models are pretrained on one or two hospital sites and are limited in scale to around 100M parameters [75; 97]. Their advantage over feature-based models in cancer risk prediction remains unclear, for instance, while CLMBR performs well on procedure outcomes and lab predictions, it fails to outperform feature-based models in pancreatic cancer prediction on EHRSHOT [97].
3 NHIRD-Cancer Benchmark
Initially, we overview our source data, NHIRD, and then the curation process of NHIRD-Cancer.
3.1 NHIRD Preliminary
Overview. The National Health Insurance Research Database (NHIRD) include electronic health record (EHR) of over 99.99% of Taiwan population. It includes decades of de-identified records from all healthcare encounters under the National Health Insurance program, with diagnoses, prescriptions, and procedures [35]. Coding standards are applied uniformly across providers, ensuring consistency and correctness. A sample patient record is shown in Appendix Figure 6(b).
Availability. The sensitivity of longitudinal EHR data makes large-scale public access challenging [29]. Among existing longitudinal EHRs [47; 75; 97], NHIRD is among the most accessible for open research. Eligible institutions in Taiwan and their international collaborators can access NHIRD with IRB approval or exemption under standardized protocols. Access is regulated by the Ministry of Health and Welfare to ensure strong privacy protections [35; 53; 82]. See Appendices A and B for details on de-identification, data quality, and access to NHIRD.
| Item | Count |
| of patients | 3,989,369 |
| of visits (In/out-patients, pharmacy) | 1,441,453,071 |
| of diagnosis (ICD9 codes) | 1,491,556,480 |
| of procedures (Surgeries) | 46,282,990 |
| of treatments (Non drug treatments) | 1,729,212,760 |
| of medications (Drug prescriptions) | 2,770,631,039 |
| of unique medical codes | 185,138 |
| of tokens for all codes | 7,923,387,479 |
| of patients for pretraining (80%) | 3,191,495 |
| of patients for finetuning (20%) | 797,874 |
| Avg. of visits per patient | 271 |
| Pancreatic | Liver | Lung | |
| Total | 315,827 | 294,516 | 310,205 |
| Training size (80%) | 252,662 | 235,613 | 248,164 |
| Validation size (10%) | 31,583 | 29,452 | 31,020 |
| Test size (10%) | 31,582 | 29,451 | 31,021 |
| Positive | 12,015 | 10,991 | 8,675 |
| Negative | 303,812 | 283,525 | 301,530 |
| Avg. history | 12.3 years | 13.8 years | 14 years |
3.2 NHIRD-Cancer Benchmark Curation
Our cancer screening benchmark is curated collaboratively with clinicians, following their guidance in the selection of positive group and cohort control-case group.
Data Statistics. We accessed a subset of NHIRD via collaboration with a Taiwanese medical school, with IRB exemption, as the study uses anonymized, non-human-subjects data. The dataset includes 3+ million patients from 1996–2013, covering 1.4 billion visits with diagnoses, treatments, and medications coded in ICD-9 [8]. It provides 15+ years of physician-confirmed, government-validated history and over 8 billion medical codes. Summary statistics are in Table 1, with preprocessing and additional details in Appendix A. Our subset aligns with full NHIRD averages (18 vs. 16.5 visits per patient over 15 years) and is comparable in scale to the 3.67M-patient records from Stanford medicine used in CLMBR [97], which is not publicly available.
Targeted Cancers. We focus on pancreatic, liver, and lung cancers, which have high mortality rates and increased risk of secondary cancer [21; 30; 39] where early detection offers substantial benefit [48; 85; 5; 32]. The screening task is a binary classification: predicting whether a patient will later develop the target cancer. As suggested by clinicians, we focus on predicting cancer one year after a clinical visit, enabling early intervention for proactive care. We set the available medical history per patient to be five years, following clinical standards.
Case and Control Group. Guided by clinicians, we rigorously follow the standard case-control study methodology and previous cancer screening studies on NHIRD [63; 49; 50] to create case-control dataset. We use the first occurance of target cancer type as the positive case, but note that patients may have other types of cancers, which we consider as the reoccurred group. The other patients are first cancer patients. Screening for both is evaluated. The (negative) control group includes patients with no cancer diagnoses, ensuring a clear distinction between cases and controls, and are matched with the case group as follows.
1.
Patient Demographic Matching: Control patients are matched to case patients based on age and gender, to minimize confounding factors from demographics.
2.
Relative Duration Matching: Control patients are then filtered to have clinical visits on the same index date as the case patient’s diagnosis date, ensuring aligned medical timelines. They are also matched with a comparable length of medical history up to the index date.
3.
Cumulative Duration Matching: Finally, controls are matched to case patients with the same total duration of lifetime medical history.
Our case-control matching goes beyond standard demographics by incorporating factors such as history-length and timing. This approach ensures closer alignment between case and control patients, accounting for temporal and historical factors, ensuring the observed differences reflect the presence or absence of cancer, thereby enhancing reliability and validity.
Dataset Processing. We use three-digit ICD-9 codes 157, 155, and 162 to define pancreatic, liver, and lung cancers, respectively. Control patients must not have cancer-related codes from 140–239 (top level of Neoplasms). To support broader cancer types and maximize early detection [13], we use top-level ICD codes to capture more cases. Table 2 lists benchmark statistics. All cancer diagnoses are confirmed by licensed physicians and validated by the government. More details in Appendix A.
Availability. NHIRD-Cancer follows the same access control as NHIRD. Approved NHIRD users can request a copy through a lightweight approval process. All data processing and benchmark curation code will be open-sourced for use with other EHR datasets.
4 Methods
This section presents the model, training, and inference of CATCH-FM.
4.1 Model
Figure 1 illustrates the model setup of CATCH-FM. It represents the healthcare record of each patient as a medical code sequence and models them by a decoder-only transformer.
Patient Representation. In longitudinal structured EHR databases, the patient records are often represented by a coding system, for example, ICD [8] or SNOMED [19], using a unique ID to encode each specific piece of information. A typical patient record can start with their demographic information, such as age and gender , followed by a sequence of chronologically ordered visits, . Each visit consists of medical events, , covering diagnoses, medications, procedures, and other clinical events.
Specifically, CATCH-FM represents each patient’s EHR token sequence as,
The demographic tokens age is discretized into predefined categorical ranges and gender is assigned as a distinct token. Each medical event code is assigned a unique token. Time intervals between consecutive visits are captured using time tokens , which are also discretized into predefined categories to encode temporal information. A special token, [EOS] marks the end of the record.
Though each ID can be mapped into the language space, for example, to their names and descriptions, CATCH-FM directly operates in the ID space which is more compact—each event is encoded by one token—and precise without ambiguity.
Architecture. CATCH-FM uses the standard decoder-only transformer architecture on top of the patient’s EHR token sequence . We use rotary positional embeddings (RoPE) [78] to encode positional information for visits:
It assigns the same absolute position (i.e., replacing sequential positions like 0,1,2,3… with 0,0,1,1…) to event tokens from the same visit. This allows the model to capture the relationships between events within and across visits.
4.2 Pretraining, Finetuning, and Inference
CATCH-FM employs the standard pretraining, Finetuning, and inference pipeline.
Pretraining. We pretrain CATCH-FM from scratch on our healthcare record pretraining dataset, using the standard autoregressive next-token prediction objective:
The next code probability is computed on token embeddings and hidden states from . Pretraining on large scale patient record with autoregressive language modeling task enables CATCH-FM to capture the complex medical patterns in patient health trajectories, for example, associations of medical events and potential risk factors of diseases.
Finetuning. For the cancer prediction task, we employs supervised fine-tuning on the pretrained CATCH-FM . It uses cross-entropy loss to learn whether a patient will be diagnosed as a target cancer outcome .
The learnable parameters include in the foundation model, and and from a linear prediction layer. Finetuning makes CATCH-FM specialized for cancer risk prediction by capturing risk factors from medical histories while leveraging general medical knowledge learned from pretraining.
Inference. The inference of CATCH-FM is to take a patient’s medical history from their EHR record, run a forward pass of the pretrained and finetuned CATCH-FM instances, and predict the cancer risk of the patient. The inference is efficient as only one prediction token is needed per patient.
The sole requirement of patient history makes CATCH-FM a nature fit for cancer prescreening. Healthcare providers can deploy CATCH-FM on a large amount of patient EHR data. The predicted high-risk patients can then be further evaluated by professionals to determine who and when to undergo cancer screening. An accurate prescreening thus effectively triages patients, improving the efficiency and efficacy of cancer screening. It can potentially increase the cancer screening rate by providing decision evidences for healthcare professionals and raise awareness from patients.
5 Experimental Methodologies
Dataset. We use the NHIRD-Cancer benchmark described in Sec. 3.2 for the pretraining, finetuning, and in-domain evaluation. There is no patient overlap among the three. To evaluate model generalization, we use the EHRSHOT benchmark [97], containing 6,739 Stanford Medicine patients (1990–2023), primarily coded in SNOMED, CPT, and RxNorm. It introduces significant distribution shifts from NHIRD: different healthcare systems, different medical codings, and different populations with different cancer prevalence. We use EHRSHOT’s official pancreatic cancer risk prediction task following their few-shot setting (1–128 positives). It is the only available cancer risk prediction task.
As CATCH-FM is pretrained on NHIRD using ICD-9 and NHI codes, we map SNOMED to ICD-9 using official CMS and NLM mappings. While 38% of codes have one-to-one mappings (mostly SNOMED to ICD-9), the remaining 62% (e.g., RxNorm, CPT) are aligned to NHI codes via semantic matching on text descriptions [56; 81; 104]. Due to differences in drug, procedure, and medical codes, the average cosine similarity between EHRSHOT and top-1 matched NHIRD codes is 84.3%, posing a significant generalization challenge. Appendix C details the mapping process.
Evaluation metrics. NHIRD-Cancer evaluations use AUPRC, specificity, sensitivity, and Negative Predictive Value (NPV). AUPRC captures the precision-recall trade-off in imbalanced tasks like cancer risk. Specificity, the ratio of true negatives to all negatives, reflects reliability in identifying low-risk patients. NPV measures confidence in negative results and is often a key metric for screening [28]. Sensitivity, the ratio of true positives to total positives, reflects how well the model flags cancer-risk patients. When possible, we evaluate sensitivity at 99% specificity, a common (pre)screening threshold [7], and a clinician-defined standard to reduce false positives, ensuring clinical trust. EHRSHOT uses AUPRC and AUROC as official metrics, and we follow exact setups.
Baselines. We include standard tree-based models with bag-of-words as input [3]: XGBoost [12] and LightGBM [45], and well-established EHR Deep learning models [92]: StageNet [26] and RETAIN [15]. We compare with pretrained language models, by converting medical codes into text and finetuning BioGPT [58] and long input sequence Qwen2.5 [100] on the same data. For EHRSHOT, we directly compare CATCH-FM against methods reported in their leaderboard [97].
Implementation Details. We implement CATCH-FM using a standard architecture [4], a decoder-only transformer with rotary embeddings and flash attention, and pretrain from scratch on NHIRD. Medical codes, demographics, and special tokens form a 185,138 token vocabulary, with sequences capped at 2,048 tokens. Longer histories are chunked without overlap during pretraining and truncated for fine-tuning. More architecture, hyperparameters, and training details are provided in Appendix E.
6 Evaluation Results
This section presents experimental results evaluating CATCH-FM’s scaling law, effectiveness, generalization, and ablation studies, as well as analyses of its captured cancer risk factors.
| FLOPs | Parameters | A100 Hours |
| 1e18 | 160m | 16 |
| 4e18 | 410m | 64 |
| 2e19 | 720m | 128 |
| 4e19 | 1b | 256 |
| 2e20 | 2.4b | 1280 |
6.1 Scaling Laws of Pretraining on Healthcare Records
To pretrain effectively on healthcare records, we first conduct a thorough scaling law analysis of CATCH-FM. Specifically, we use IsoFLOP profiling [34]: pretraining models with various sizes at target FLOPs by varying the number of pretraining tokens.
Figure 2(a) illustrates clear parabola-shaped IsoFLOP curves, with different model sizes achieving minimum validation loss at various FLOPs. Using these data points, we fit a power law to characterize the relationship between FLOPs , the loss-optimal model size , and the optimal number of training tokens , as illustrated in Figures 2(b) and 2(c):
Scaling laws in healthcare foundation models resemble the scaling laws of large language models [34], highlighting the potential of large-scale EHR foundation models. While emergent capabilities need further study, we pretrain and finetune a series of compute-optimal CATCH-FM models across various FLOPs (Table 3) and evaluate their efficacy in cancer risk prediction. The benefits of compute-optimal pretraining are detailed in Appendix G.
| Model | Pancreatic (1201/30382, 3.95%) | Liver (1099/28353, 3.88%) | Lung (868/30153, 2.88%) | ||||||
| AUPRC | Speci. NPV. | Sensi. | AUPRC | Speci. NPV. | Sensi. | AUPRC | Speci. NPV. | Sensi. | |
| XGBoost | 68.8 | 99.6 97.8 | 46.6 | 69.2 | 99.6 97.9 | 47.1 | 64.5 | 99.7 98.3 | 42.3 |
| LightGBM | 72.9 | 99.6 98.0 | 51.3 | 73.1 | 99.6 97.8 | 52.3 | 68.9 | 99.7 98.4 | 47.0 |
| RETAIN | 11.6 | 80.7 97.3 | 46.0 | 8.8 | 85.4 97.4 | 44.3 | 2.1 | 100.0 97.1 | 0.0 |
| StageNet | 11.8 | 100.0 96.0 | 0.0 | 9.0 | 100.0 96.1 | 0.0 | 6.2 | 100.0 97.1 | 0.0 |
| BioGPT-347m | 55.0 | 98.8 97.7 | 44.1 | 55.4 | 99.3 97.5 | 38.4 | 58.5 | 99.4 98.3 | 44.5 |
| Qwen2.5-500m | 73.7 | 98.9 98.4 | 62.4 | 70.0 | 99.4 98.2 | 55.0 | 70.2 | 99.3 98.7 | 57.1 |
| CATCH-FM-160m | 81.2 | 99.2 98.7 | 70.7 | 76.6 | 99.3 98.5 | 63.4 | 71.5 | 99.6 98.6 | 54.7 |
| CATCH-FM-410m | 81.2 | 99.3 98.7 | 70.8 | 78.5 | 99.5 98.5 | 63.6 | 73.6 | 99.7 98.6 | 55.3 |
| CATCH-FM-720m | 82.0 | 99.2 98.8 | 73.1 | 76.5 | 99.3 98.6 | 66.1 | 69.1 | 99.2 98.8 | 61.9 |
| CATCH-FM-1b | 82.7 | 99.2 98.9 | 73.5 | 77.0 | 99.1 98.7 | 68.7 | 69.6 | 99.2 98.8 | 62.0 |
| CATCH-FM-2.4b | 82.5 | 99.2 98.9 | 73.7 | 76.9 | 99.1 98.7 | 69.0 | 69.6 | 99.3 98.9 | 63.3 |
6.2 Overall Results
Table 4 shows the overall performance of cancer risk prediction in NHIRD-Cancer. Cancer risk prediction is a challenging task. Many baselines fail to achieve meaningful performance. CATCH-FM is the only method that achieves 60%+ sensitivity on all three types of cancer while maintaining nearly 99% specificity and NPV, confirming its ability to identify high-risk patients for further screening (high sensitivity) and avoid unnecessary patient distress (high specificity).
CATCH-FM outperforms XGBoost and LightGBM, strong tree models that often outperformed EHR foundation models [97], by 10+% sensitivity. It also significantly outperforms medical (BioGPT) and general (Qwen) large language models pretrained on the texts of NHIRD. Sect. 6.4 further studies the benefits of operating on medical codes. Note that the specificity-sensitivity trade-off can be easily adjusted for different scenarios by changing classification thresholds. Our studies show that CATCH-FM maintains its advantage across all thresholds. Additional details are found in Appendix F.
| Model | Pancreatic | Liver | Lung | ||||||
| AUPRC | Speci. NPV. | Sensi. | AUPRC | Speci. NPV. | Sensi. | AUPRC | Speci. NPV. | Sensi. | |
| Training Cohorts: all cancer patients, Testing Cohorts: first cancer patients | |||||||||
| Pos/Neg: | (559/30382 1.83%) | (528/28353 1.86%) | (419/30153 1.39%) | ||||||
| XGBoost | 59.2 | 99.8 98.8 | 38.6 | 65.2 | 99.8 98.9 | 44.1 | 61.8 | 99.9 99.1 | 41.8 |
| Qwen2.5-500m | 66.6 | 98.9 99.2 | 61.7 | 64.4 | 99.4 99.1 | 55.9 | 62.5 | 99.3 99.3 | 56.1 |
| CATCH-FM-160m | 56.8 | 99.3 99.1 | 52.5 | 48.9 | 99.3 98.9 | 43.9 | 40.7 | 99.6 99.1 | 34.9 |
| CATCH-FM-1b | 60.3 | 99.2 99.1 | 55.4 | 50.0 | 99.1 99.1 | 50.6 | 40.7 | 99.2 99.2 | 41.1 |
| CATCH-FM-2.4b | 60.0 | 99.2 99.2 | 57.6 | 51.4 | 99.1 99.1 | 51.7 | 41.4 | 99.3 99.2 | 45.5 |
| Training Cohorts: all cancer patients, Testing Cohorts: reoccurred cancer patients | |||||||||
| Pos/Neg: | (642/30382 2.11%) | (571/28353 2.01%) | (449/30153 1.49%) | ||||||
| XGBoost | 90.7 | 99.8 99.5 | 80.4 | 91.7 | 99.8 99.5 | 80.4 | 89.0 | 99.9 99.6 | 76.6 |
| Qwen2.5-500m | 67.2 | 98.9 99.2 | 63.4 | 61.4 | 99.4 99.1 | 54.0 | 63.7 | 99.3 99.3 | 53.2 |
| CATCH-FM-160m | 89.3 | 99.3 99.7 | 86.1 | 85.9 | 99.3 99.6 | 81.4 | 80.7 | 99.6 99.6 | 73.9 |
| CATCH-FM-1b | 91.4 | 99.2 99.8 | 89.3 | 86.9 | 99.1 99.7 | 85.1 | 80.5 | 99.3 99.7 | 79.0 |
| CATCH-FM-2.4b | 86.0 | 99.2 99.7 | 87.7 | 84.9 | 99.1 99.7 | 85.0 | 78.8 | 99.3 99.7 | 80.0 |
| Training Cohorts:first cancer patients, Testing Cohorts: first cancer patients | |||||||||
| Pos/Neg: | (452/28058 1.59%) | (509/26148 1.95%) | (399/27396 1.45%) | ||||||
| XGBoost | 26.3 | 99.9 99.5 | 10.2 | 36.2 | 99.8 98.4 | 22.6 | 27.6 | 99.9 98.7 | 10.3 |
| Qwen2.5-500m | 22.0 | 99.3 98.6 | 17.9 | 32.4 | 99.0 98.5 | 29.2 | 15.9 | 99.1 98.8 | 18.4 |
| CATCH-FM-160m | 42.4 | 99.4 98.9 | 37.4 | 39.3 | 99.4 98.7 | 33.8 | 33.0 | 99.5 99.0 | 29.6 |
| CATCH-FM-1b | 57.2 | 99.6 99.2 | 48.7 | 49.6 | 99.4 98.9 | 44.4 | 47.5 | 99.4 99.2 | 45.6 |
| CATCH-FM-2.4b | 61.3 | 99.6 99.2 | 53.4 | 52.8 | 99.2 99.0 | 49.3 | 49.6 | 99.4 99.2 | 47.7 |
Table 5 shows CATCH-FM’s performances when finetuned with all cancer patients or only with the first cancer patients. Over half of the target patients fall into the reoccurred group, where CATCH-FM achieves 80% sensitivity, effectively leveraging previous cancer diagnostics as strong risk indicators. Detecting first cases is much more challenging. While other methods fail to balance the effectiveness of the two groups, CATCH-FM achieves strong effectiveness on both first and reoccurred patients when finetuned with the two groups together.
The benefit of scale is most evident in predicting the risks of the first cancer patients. CATCH-FM-2.4b outperforms CATCH-FM-160m by 5–10% sensitivity when finetuned on all cancer patients, and by more than 15% when trained exclusively in first-time cancer cases. Scaled-up foundation models perform better on harder tasks and with fewer finetuning labels. We further examine CATCH-FM’s generalization in the next experiment. Results with limited labels are in Appendix M.
6.3 Generalization Ability
Across Hospital Site. Figure 4 shows the K-shot results of CATCH-FM on the EHRSHOT official pancreatic cancer leaderboard, using K positive and K negative on-site examples for tuning. Despite significant shifts in population, healthcare systems, and noise in medical code mapping, CATCH-FM achieves state-of-the-art AUPRC across all shots on the EHRSHOT leaderboard with a handful of on-site examples, maintains the best AUROC for K64, and ranks second only at K=128 with a marginal difference (<3%). Moreover, CATCH-FM-160m, at a similar scale, outperforms CLMBR (140m) in AUPRC across all shots. Although its AUROC is lower, scaling to CATCH-FM-2.4b significantly improves few-shot performance, showing the benefits of scale. Notably, CLMBR is pretrained at a similar scale but with on-site data: 2.57M patient records from Stanford Medicine, the same as EHRSHOT, while CATCH-FM is challenged by distribution shifts and noisy medical code mappings. These results demonstrate CATCH-FM’s potential to transfer across different hospital sites.
Across Population. This experiment evaluates CATCH-FM on cohorts when cancer prescreening targets the general population with much lower cancer prevalence [71], and on randomly sampled controls. Figure 4 plots the performance with varying positive ratio and shows CATCH-FM’s robustness to different population distributions. For those interested, Appendix includes additional experiments on CATCH-FM’s robustness across other cohort variations (H), different history exclusion window (I), demographics (J), and its performance on other diagnostic support tasks (K).
| Variations | AUPRC | Specificity | Sensitivity |
| Medical Code Representation | |||
| CATCH-FM | 81.2 | 99.2 | 70.7 |
| demographic-only features | 51.9 | 100 | 0 |
| No Pretrain | 71.2 | 99.3 | 55.7 |
| Model Architecture | |||
| w. Token Level Rel-Pos | 81.1 | 99.2 | 70.4 |
| w/o. time token | 80.0 | 99.2 | 70.3 |
| Converted Text Representation | |||
| demographic-only features | 52.9 | 100 | 0 |
| No Pretrain | 39.9 | 96.5 | 49.6 |
| Finetune Pythia | 59.4 | 98.7 | 51.8 |
| Pretrain from Scratch | 64.8 | 98.9 | 52.6 |
| Continual Pretrain Pythia | 64.6 | 98.2 | 57.3 |
6.4 Ablation Study
Table 6 shows ablation studies on different components of CATCH-FM. Modeling patient history is a key source of evidence for CATCH-FM. Demographic information alone is a poor indicator; pretraining is a key advantage of CATCH-FM; different model architectures, similar to observations in large language models, yield mild differences. More detailed ablations are presented in Appendix N
Converting the medical codes into their corresponding textual names and pretraining CATCH-FM on texts hinders performance. As shown in Figure 5(a), the loss quickly drops to near zero, a strong sign of overfitting. The model memorizes code names and predicts them trivially after seeing the initial tokens. Figure 5(b) shows the loss is only meaningful on the first token, with the rest memorized. How to better adapt general domain LLMs to EHR data is an interesting future research direction.
| Cancer | Features Learned by Sparse Autoencoder [27] with Description Generated by Node2Graph [25] |
| Pancreatic | High blood pressure, Diseases of the respiratory system, Operations on larynx or trachea, Telmisartan |
| Type 2 diabetes mellitus, Hypertensive heart disease, Norvasc tablet, Hypothyroidism and thyroid | |
| Infectious and parasitic diseases, Intestinal infectious diseases, Pramipexole, Piracetam | |
| Liver | Duodenal ulcer, Tetracyclines, Ursodeoxycholic acid, Famotidine |
| Repair of uterus and supporting structures, Diagnosis on lymphatic structures, Ancillin, Ketoprofen | |
| Diclofenac, Anxiety state, Dysthymic disorder, Gastrojejunal ulcer | |
| Lung | Loperamide, Chest view, Chronic ischemic heart disease, Diovan, Lymphatic diagnostic procedures |
| Hypertensive disease with congestive heart failure, Isosorbide mononitrate, Inguinal hernia repair | |
| Adenocarcinoma, Esophagomyotomy, Norvasc tablet, Isosorbide mononitrate |
6.5 Interpretability Analysis
This experiment leverages recently developed LLM interpretation method [27; 44] to understand the risk factors captured by CATCH-FM. We train a sparse autoencoder on the prediction layers of finetuned CATCH-FM and then leverage neuron-to-Graph (N2G) [25] to explain the top active latent features on positive cancer cases. Additional implementation details are found in Appendix L. Table 7 identifies features corresponding to top cancer risk factors captured by CATCH-FM. It includes not only trivial factors, such as Type 2 diabetes [20] and high blood pressure [77], but also non-trivial ones such as hypothyroidism, thyroid disorders, and Norvasc, with some indicators only presented at top medical journals less than ten years ago [69; 95], after our training data cutoff.
7 Conclusion
CATCH-FM provides a new cancer risk prediction methodology by pretraining and finetuning Transformers on millions of patients’ medical code sequences. Its effectiveness (60%+ sensitivity), low risk (99% specificity), and wide applicability (only requiring inference on EHR records) make it a natural fit for cancer pre-screening. It helps healthcare professionals efficiently decide whom and when to screen for cancer, potentially improving the effectiveness and coverage of cancer screening and ultimately, patient outcomes. Our experiments on NHIRD and EHRSHOT demonstrated the benefit of scale in pretraining EHR foundation models, their generalization ability across significantly different healthcare systems, and their ability to capture non-trivial cancer risk factors. We hope our findings, analyses, and open-source codes can inspire and facilitate further research and deployments in leveraging AI to solve real-world healthcare problems.
Acknowledgments
We sincerely thank CMU Foundation and Language Model (FLAME) Center for providing support of computational resources. We also thank Fu Jen Catholic University and Hospital for providing data access and clinical collaboration support.
Appendix A NHIRD Details
The Taiwanese National Health Insurance Research Database (NHIRD) is one of the largest and most comprehensive de-identified population-wide EHR datasets globally [35], covering over 99% of the Taiwanese population. It contains longitudinal medical records spanning over two decades, including patient demographics, diagnoses, prescriptions, clinical events, medical procedures, and hospital visits from all hospitals and medical facilities in Taiwan. Maintained by the National Health Insurance Administration in collaboration with the Ministry of Health and Welfare, the NHIRD comprises both registration files and hospital claim data submitted for reimbursement under the National Health Insurance (NHI) program.
We leverage a subset, which includes a randomly selected cohort of three million patients from 1996 to 2013, of the Taiwanese National Health Insurance Research Database (NHIRD) for pretraining EHR foundation models and constructing clinical downstream tasks for benchmarks. Unlike other datasets, the NHIRD stands out for its scale and comprehensiveness under a single-payer healthcare system, which enables standardized data collection and creates a comprehensive and lifelong record of patients’ medical footprints. This makes NHIRD one of the most suitable resources for the real-world implementation of EHR foundation models. The NHIRD contains three main categories of medical information, with all personal details, such as ID, birthdate, and residential postcode, de-identified:
Demographics: This includes details about medical institutions (e.g., centers, hospitals, and clinics) and de-identified data on patients, physicians, and pharmacists.
Visit: Comprehensive records of medical visits, including outpatient clinic visits (including visit to medical centers and hospitals), hospitalizations, and pharmacy drug fulfillments.
Order details: Detailed records of prescriptions, procedures, medical equipment, and materials associated with each type of visit.
Table 8 lists all NHIRD tables with descriptions. The three main visit tables are CD (outpatient clinic visits, including medical centers and hospitals), DD (hospitalization), and GD (pharmacy), each linked to an order table: OO (for CD), DO (for DD), and GO (for GD). For CATCH-FM, we use ID, CD, OO, DD, DO, GD, and GO tables, as they constitute complete patient medical histories. Demographic and other statistics to further describe data density and sparsity are shown in Table 9 and Figure 7. Rigorous data cleaning includes removing missing values, correcting erroneous codes, and standardizing dates and billing codes to ensure reliable pretraining data. The following sections detail how we construct patient medical histories and preprocess data to create the cancer screening benchmark.
Data De-identification and Privacy. All NHIRD data are de-identified following strict protocols mandated by the National Health Insurance Administration (NHIA). Personally identifiable information (PII), such as names and ID numbers, is removed and irreversibly encrypted using non-public anonymization methods [53; 33]. Our study uses only anonymized demographic variables, e.g., age and gender, for model training. Given the level of anonymization and the coarse granularity of these attributes, the risk of patient re-identification is negligible. Hence, our use of NHIRD data fully complies with privacy standards and poses no ethical or legal concerns regarding patient confidentiality.
Correctness of NHIRD. The universal healthcare system in Taiwan enables frequent patient visits and interactions and generates rich, detailed, and longitudinal patient records. In 2023, NHIRD recorded over 380 million medical visits for 23 million individuals, averaging 16.5 visits per person [1]. These statistics, including consistent surgery and prescription rates, align closely with our dataset statistics in Table 1, confirming the correctness of our dataset. Most importantly, all data in NHIRD receive strict validation by the National Health Insurance Administration (NHIA), which enforces robust quality control measures to eliminate duplication, correct inconsistencies, and ensure patient-level accuracy. Additionally, coding standards are uniformly implemented across all healthcare providers nationwide, guaranteeing consistency and correctness at scale. Given this systematic validation, national standardization, and population-wide coverage, NHIRD is an exceptionally accurate and dependable data source trusted and used in research published at top peer-reviewed medical journals [93; 51; 89; 83].
Cancer Diagnosis Validity. Cancer diagnoses in the NHIRD are made by licensed physicians from all hospitals and medical facilities and undergo strict validation by the National Health Insurance Administration (NHIA) to ensure diagnostic accuracy, prevent misclassification, and eliminate billing errors or fraud [53; 35]. This rigorous quality control process makes the NHIRD a highly reliable source for cancer retrospective research, especially for creating cancer patient cohorts with diagnosis codes. With verified patient histories and clinically validated cancer diagnoses, the NHIRD has been widely adopted for cancer research and trusted in top peer-reviewed studies [54; 14; 37].
| Name | Description | Name | Description | ||
| BED | Registry for contracted beds | ID* |
| DETA | Registry for contracted specialty services | DT | Monthly claim summary for inpatient claims | ||
| HOSB | Registry for contracted medical facilities | CT | Monthly claim summary for ambulatory care claims | ||
| HOSX | Supplementary registry for contracted medical facilities | DD* |
| DOC | Registry for board-certified specialists | DO* |
| PER | Registry for medical personnel | CD* |
| HV | Registry for catastrophic illness patients | OO* |
| HOX | Registry for medical services | GD* |
| DRUG | Registry for drug prescriptions | GO* |
| Group | Count | Avg. history | Median history | Avg. # visits | Median # visits | Avg. # codes | Median # codes |
| All | 3,989,369 | 15.2 | 17 | 271 | 214 | 5,886 | 4,329 |
| Male | 1,965,368 | 15.1 | 17 | 243 | 184 | 5,415 | 3,818 |
| Female | 1,962,523 | 16 | 17 | 306 | 248 | 6,504 | 4,934 |
| 0-18 | 456,322 | 12.4 | 13 | 274 | 250 | 5,803 | 5,208 |
| 18-35 | 1,047,623 | 15.4 | 17 | 188 | 164 | 3,736 | 3,137 |
| 35-50 | 981,514 | 15.5 | 17 | 214 | 172 | 4,448 | 3,384 |
| 50-70 | 1,017,522 | 16 | 17 | 312 | 253 | 6,829 | 5,246 |
| 70+ | 486,388 | 15 | 17 | 477 | 416 | 11,525 | 9,871 |
Medical History Construction. To create a sequential medical history for each patient, we first aggregate all visits and their associated order details by joining CD with OO, DD with DO, and GD with GO. Figure 6(a) describes the join process on CD with OO. After that, we aggregate visits for each individual 4 million patients by their patient ID and sorted in chronological order. Figure 6(b) shows a sample patient after the aggregation on all visits.
Cancer Demographic Statistics. Table 10 presents demographic statistics, including cancer patients’ medical history lengths across age groups. The NHIRD provides extensive longitudinal data, with average history lengths ranging from 10.3 to 15 years across various cancers. This depth allows for analyzing disease progression and identifying patterns that support early cancer detection. Its large sample size ensures robust statistical power for subgroup analysis, making NHIRD an invaluable resource for population-wide studies on cancer progression, early detection, and screening effectiveness.
| Demographic group | Counts | Avg. history length |
| Pancreatic | ||
| All | 12,602 | 12.3 |
| 0-18 | 28 | 11.8 |
| 18-35 | 267 | 15 |
| 35-50 | 1,453 | 13.3 |
| 50-70 | 4,806 | 12.8 |
| 70+ | 6,048 | 11.6 |
| Liver | ||
| All | 11,355 | 13.7 |
| 0-18 | 14 | 10.3 |
| 18-35 | 252 | 14.3 |
| 35-50 | 1,620 | 14.2 |
| 50-70 | 4,968 | 13.9 |
| 70+ | 4,501 | 13.1 |
| Lung | ||
| All | 8,867 | 14 |
| 0-18 | 15 | 12.8 |
| 18-35 | 120 | 14.5 |
| 35-50 | 678 | 14.3 |
| 50-70 | 3,109 | 14.4 |
| 70+ | 4,945 | 13.8 |
Appendix B Data Accessibility, IRB, and Reproducibility
NHIRD is a publicly accessible research resource governed under controlled access. Any research institution in Taiwan may apply for access by obtaining Institutional Review Board (IRB) approval and complying with the data use regulations set by the Ministry of Health and Welfare. International research institutions can collaborate with Taiwanese research institutions under a formal agreement. While the data are not openly accessible, they are available to all eligible institutions through a formal application process. We obtained NHIRD access through a formal collaboration with a Taiwanese medical school. As part of the IRB review process, we submitted the complete NHIRD data schema, a comprehensive research plan, the data access protocol, and a list of authorized users. We received formal approval and granted exemption from our IRB, as the study relied solely on fully de-identified secondary data and involved no human subjects. Such IRB determination and exemption are common in research using the MIMIC dataset. For reproducibility, we will publicly release our complete data preprocessing pipeline and modeling code. In addition, pre-trained model checkpoints are made available to researchers who obtain NHIRD access through the same framework, enabling fair replication and downstream research while upholding strict privacy and compliance standards.
Appendix C EHRSHOT
Dataset and Prediction Task Overview. EHRSHOT is the latest published EHR dataset for evaluating the few-shot performance of foundation models on clinical prediction tasks. It consists of structured, de-identified, and longitudinal Electronic Health Records (EHRs) from 6,739 patients treated at Stanford Medicine, including diagnosis codes, procedures, medications, and laboratory test results. EHRSHOT is sourced from the EHR system in both Stanford Health Care and Lucile Packard Children’s Hospital, where 2.57M patients used to pretrain CLMBR, a clinical foundation model used to evaluate the EHRSHOT benchmark in the original paper [97]. It contains 15 distinct clinical prediction tasks under the four main categories:
Operational Outcomes
Anticipating Lab Test Results
Assignment of New Diagnoses
Anticipating Chest X-ray Findings
where the one-year pancreatic cancer prediction is under the category "Assignment of New Diagnoses". It is worth noting that, pancreatic cancer prediction is the only available cancer prediction task in EHRSHOT. Each designed to test the ability of foundation models for accurate predictions under limited labeled data at the time of patient visit. Refer Table 3 in the original EHRSHOT paper [97] for complete task description and statistics.
| Code type | EHRSHOT | NHIRD |
| Disgnosis | SNOMED, ICDO3 | ICD-9-CM |
| Prescriptions | RxNorm | NHI Drug Code |
| Medical Procedures | CPT, HCPCS | NHI Order Code |
| Surgical Procedures | SNOMED, ICD-10-Proc, ICD-9-Proc, | ICD-9-Proc, NHI Order Code |
| Diagnositc Procedures | CPT, HCPCS | NHI Order Code |
| Lab Orders | LONIC | NHI Order Code |
| Medical Equipments | SNOMED, CPT, HCPCS | NHI Material Code |
| Medical Supplies | HCPCS | NHI Material Code |
| Medical Services | CPT | NHI Service Code |
EHRSHOT verse NHIRD. The source EHR system of EHRSHOT in Stanford Medicine and Lucile Packard Children’s Hospital utilizes the Observational Medical Outcomes Partnership Common Data Model (OMOP-CDM) format. The OMOP-CDM is a standardized format for organizing and conforming EHR, enabling consistent analysis across diverse healthcare datasets. On the other hand, The NHIRD, as described in Appendix A, comprises an extensive set of structured tables capturing nearly all aspects of healthcare encounters. However, its format is tailored for internal use by the National Health Insurance Administration in Taiwan, rather than for international interoperability like OMOP-CDM. Applying CATCH-FM, built and pretrained on NHIRD, to EHRSHOT is nontrivial due to differences in coding formats between the two healthcare systems. Table 11 summarizes the key differences in medical coding schemes used by NHIRD and EHRSHOT.
| Code in EHRSHOT | Mapping methods | Code in NHIRD |
| SNOMED | SNOMED to ICD9 mapping | ICD-9-CM |
| SNOMED - procedure | Semantic text matching | ICD-9-Proc, NHI Order Code |
| SNOMED - regime/therapy | Semantic text matching | ICD-9-Proc, NHI Order Code |
| SNOMED - physical object | Semantic text matching | NHI Order Code |
| ICDO3 | ICD mapping by CMS | ICD-9-CM |
| ICD10-Proc | ICD mapping by CMS | ICD-9-Proc |
| RxNorm | Semantic text matching | NHI Drug Code |
| CPT | Semantic text matching | NHI Order / Service Code |
| HCPCS | Semantic text matching | NHI Material Code |
| LONIC | Semantic text matching | NHI Order Code |
Mapping EHRSHOT to NHIRD. ICD-based codes such as ICD-9-Proc (ICD-9 Procedure code), ICD-10-PCS (ICD-10 Procedure code), and ICD-O-3 (ICD oncology code) can be mapped to ICD-9 using publicly available mapping from the Centers for Medicare & Medicaid Services (CMS). Diagnostic concepts in SNOMED can also be mapped to ICD-9 via the "ICD-9-CM Diagnostic Codes to SNOMED CT Map" provided by the U.S. National Library of Medicine (NLM). However, other coding systems, e.g., SNOMED procedures, CPT (coding for medical services, procedures, and other practices), and RxNorm (codes for drug prescriptions), lack direct mappings to the NHIRD coding scheme. To address this, we adopt a text-based semantic matching approach by embedding the textual descriptions of codes using Sentence Transformers [84] and formulating the mapping as a dense retrieval problem, where several prior works adopt this method to align different medical concepts and terms [81; 56; 104]. Codes from each coding system are matched to the most semantically similar code in NHIRD based on text similarity. We use Faiss as our dense retrieval framework [22]. Table 12 summarizes the mapping strategies and source-target relationships between each coding system. With the fundamental structural and terminological differences between these two EHR systems, our average top-1 semantic alignment score (cosine similarity) between code in EHRSHOT and code in NHIRD is approximately 84.3%. This score reflects the inherent challenge of mapping between heterogeneous and non-standardized EHR codebases, especially when deploying healthcare foundation models across healthcare systems and populations, showing the difficulties of model generalization.
Appendix D NPV evaluation
Here we present the formula to calculate Negative Predictive Value (NPV) given screening model Sensitivity, Specificity, and Prevalence [59]:
where Sensitivity, Specificity, and Prevalence are in decimal, and we use positive rate for each target cancer as Prevalence.
| Parameters | Num of layers | Dimension | Num of heads | Block size |
| 70m | 6 | 512 | 8 | 2048 |
| 120m | 6 | 768 | 8 | 2048 |
| 160m | 12 | 768 | 12 | 2048 |
| 260m | 12 | 1024 | 16 | 2048 |
| 350m | 20 | 1024 | 16 | 2048 |
| 410m | 24 | 1024 | 16 | 2048 |
| 560m | 22 | 1280 | 10 | 2048 |
| 720m | 20 | 1536 | 12 | 2048 |
| 1b | 16 | 2048 | 8 | 2048 |
| 1.2b | 20 | 2048 | 16 | 2048 |
| 1.4b | 24 | 2048 | 16 | 2048 |
| 2.1b | 24 | 2560 | 16 | 2048 |
| 2.8b | 32 | 2560 | 32 | 2048 |
| Hyperparameter | Pretraining | Supervised Fine-tuning |
| Learning Rate | 6e-6 for model size 1B; else 1e-5 | |
| Optimizer | AdamW | |
| Adam | 1e-8 | |
| Adam Betas (, ) | (0.9, 0.999) | |
| Weight decay | 0.01 | |
| Gradient Norm | 0.1 | |
| Scheduler | Warmup-Stable-Decay | |
| Warmup Ratio | 0.1 | |
| Stable Ratio | 0.8 | |
| Decay Ratio | 0.1 | |
| Batch Size | 64 | 128 |
| Epochs | - | 5 |
| FLOPS | 1e+18 | 4e+18 | 8e+18 | 2e+19 | 4e+19 | 8e+19 | 1e+20 | 2e+20 |
| CATCH-FM-70m | 10071 (1.32e+09) | 40287 (5.28e+09) | 80575 (1.06e+10) | 201439 (2.64e+10) | 402878 (5.28e+10) | 805757 (1.06e+11) | 1007197 (1.32e+11) | 2014394 (2.64e+11) |
| CATCH-FM-120m | 6247 (8.19e+08) | 24990 (3.28e+09) | 49981 (6.55e+09) | 124952 (1.64e+10) | 249905 (3.28e+10) | 499810 (6.55e+10) | 624763 (8.19e+10) | 1249526 (1.64e+11) |
| CATCH-FM-160m | 4800 (6.29e+08) | 19202 (2.52e+09) | 38405 (5.03e+09) | 96014 (1.26e+10) | 192029 (2.52e+10) | 384058 (5.03e+10) | 480072 (6.29e+10) | 960145 (1.26e+11) |
| CATCH-FM-260m | 3252 (4.26e+08) | 13011 (1.71e+09) | 26022 (3.41e+09) | 65057 (8.53e+09) | 130114 (1.71e+10) | 260228 (3.41e+10) | 325285 (4.26e+10) | 650570 (8.53e+10) |
| CATCH-FM-350m | 2421 (3.17e+08) | 9685 (1.27e+09) | 19371 (2.54e+09) | 48429 (6.35e+09) | 96858 (1.27e+10) | 193716 (2.54e+10) | 242145 (3.17e+10) | 484290 (6.35e+10) |
| CATCH-FM-410m | 2147 (2.81e+08) | 8588 (1.13e+09) | 17176 (2.25e+09) | 42941 (5.63e+09) | 85882 (1.13e+10) | 171765 (2.25e+10) | 214706 (2.81e+10) | 429412 (5.63e+10) |
| CATCH-FM-720m | 1302 (1.71e+08) | 5209 (6.83e+08) | 10418 (1.37e+09) | 26045 (3.41e+09) | 52090 (6.83e+09) | 104180 (1.37e+10) | 130225 (1.71e+10) | 260451 (3.41e+10) |
| CATCH-FM-1b | 964 (1.26e+08) | 3857 (5.06e+08) | 7714 (1.01e+09) | 19285 (2.53e+09) | 38570 (5.06e+09) | 77141 (1.01e+10) | 96426 (1.26e+10) | 192853 (2.53e+10) |
| CATCH-FM-1.2b | 818 (1.07e+08) | 3273 (4.29e+08) | 6547 (8.58e+08) | 16369 (2.15e+09) | 32739 (4.29e+09) | 65478 (8.58e+09) | 81848 (1.07e+10) | 163696 (2.15e+10) |
| CATCH-FM-1.4b | 710 (9.32e+07) | 2843 (3.73e+08) | 5687 (7.46e+08) | 14219 (1.86e+09) | 28439 (3.73e+09) | 56879 (7.46e+09) | 71098 (9.32e+09) | 142197 (1.86e+10) |
| CATCH-FM-2.1b | 486 (6.38e+07) | 1946 (2.55e+08) | 3892 (5.10e+08) | 9732 (1.28e+09) | 19464 (2.55e+09) | 38929 (5.10e+09) | 48662 (6.38e+09) | 97324 (1.28e+10) |
| CATCH-FM-2.8b | 382 (5.01e+07) | 1529 (2.00e+08) | 3058 (4.01e+08) | 7646 (1.00e+09) | 15292 (2.00e+09) | 30584 (4.01e+09) | 38230 (5.01e+09) | 76460 (1.00e+10) |
Appendix E Implementation Details
Tokenization and inputs. We map all medical codes to unique indices ranging from 0 to the total number of unique medical codes, with a token vocabulary size of 185,138, including demographic, time, and special tokens. Since all tokens are treated as atomic units, no additional tokenization is required. The input sequence length is limited to 2,048 tokens. For patient records with EHR sequences exceeding this limit, we avoid truncation during pretraining and instead split the sequences into non-overlapping chunks, processing them across multiple training steps. However, during fine-tuning, inputs longer than 2,048 tokens are truncated, as they must be processed as single sequences.
Backbone architecture. We use the Pythia architecture as the backbone of CATCH-FM. Pythia is a family of decoder-only autoregressive language models, ranging from 70m to 12b parameters, designed for scalable and consistent research. Its feedforward architecture features rotary embeddings for positional encoding, untied embedding layers, and parallelized flash attention for efficient training. Table 13 outlines the model architecture, while Table 14 details the hyperparameters for pretraining and supervised fine-tuning on cancer screening tasks. All models are trained on 8 A100-SXM4-40GB GPUs.
Baselines. For the tree-based baselines, XGBoost and LightGBM, we utilize their official Python packages, xgboost 111https://xgboost.readthedocs.io/en/stable/index.html and lightgbm 222https://lightgbm.readthedocs.io/en/stable/, respectively. For deep learning baselines, we employ the PyHealth framework [101] and perform supervised learning tasks directly on the cancer screening benchmarks. For language model baselines, we convert medical codes into text using language model vocabularies and concatenate them into sequences representing patient histories for fine-tuning. In NHIRD, ICD9 codes are mapped to standardized textual descriptions using official code mapping tables. Each medical code, representing a diagnosis, surgery, treatment, or medication, is transformed accordingly. Patient histories are then constructed by concatenating the medical text, similar to processing text documents in language models. The sequence length limit is 1024 for BioGPT and 2048 for Qwen.
Training Configurations. Table 15 shows the required training steps and total tokens for CATCH-FM at various FLOP targets. We adopt this configuration to investigate the scaling laws for CATCH-FM, aiming to determine the optimal model and data scales under a fixed FLOP budget. Their corresponding FLOPs, number of parameters, and GPU training hours are listed in Table 3. We calculate the FLOPs with Pytorch built-in flops counter.
| Threshold | False Positive Rate | True Positive Rate | Specificity | Precision | Relative Risk |
| Cancer: Pancreatic, Positive/Negative (Incidence Ratio): 1201/30382 (3.95%) | |||||
| 0.999 | 0.001 | 0.600 | 0.999 | 0.956 | 25.146 |
| 0.999 | 0.001 | 0.630 | 0.999 | 0.945 | 24.853 |
| 0.997 | 0.002 | 0.654 | 0.998 | 0.932 | 24.517 |
| 0.990 | 0.003 | 0.670 | 0.997 | 0.909 | 23.893 |
| 0.975 | 0.003 | 0.684 | 0.997 | 0.886 | 23.293 |
| 0.941 | 0.004 | 0.696 | 0.996 | 0.868 | 22.826 |
| 0.825 | 0.005 | 0.707 | 0.995 | 0.836 | 21.975 |
| 0.710 | 0.007 | 0.717 | 0.993 | 0.806 | 21.220 |
| 0.550 | 0.009 | 0.728 | 0.991 | 0.780 | 20.521 |
| 0.391 | 0.010 | 0.740 | 0.990 | 0.753 | 19.796 |
| Cancer: Liver, Positive/Negative (Incidence Ratio): 1099/28353 (3.88%) | |||||
| 0.915 | 0.004 | 0.601 | 0.996 | 0.856 | 22.941 |
| 0.891 | 0.004 | 0.612 | 0.996 | 0.848 | 22.715 |
| 0.852 | 0.004 | 0.624 | 0.996 | 0.833 | 22.338 |
| 0.819 | 0.005 | 0.637 | 0.994 | 0.823 | 22.070 |
| 0.757 | 0.006 | 0.646 | 0.994 | 0.807 | 21.622 |
| 0.705 | 0.006 | 0.654 | 0.993 | 0.791 | 21.197 |
| 0.652 | 0.007 | 0.668 | 0.993 | 0.782 | 20.949 |
| 0.617 | 0.008 | 0.676 | 0.992 | 0.765 | 20.479 |
| 0.579 | 0.009 | 0.684 | 0.991 | 0.754 | 20.213 |
| 0.508 | 0.010 | 0.696 | 0.990 | 0.736 | 19.713 |
| Cancer: Lung, Positive/Negative (Incidence Ratio): 868/30153 (2.88%) | |||||
| 0.853 | 0.006 | 0.600 | 0.994 | 0.730 | 26.078 |
| 0.818 | 0.007 | 0.604 | 0.993 | 0.725 | 25.902 |
| 0.727 | 0.007 | 0.608 | 0.992 | 0.713 | 25.465 |
| 0.686 | 0.007 | 0.614 | 0.993 | 0.705 | 25.196 |
| 0.632 | 0.008 | 0.619 | 0.992 | 0.698 | 24.957 |
| 0.572 | 0.008 | 0.622 | 0.992 | 0.687 | 24.522 |
| 0.523 | 0.008 | 0.628 | 0.992 | 0.680 | 24.316 |
| 0.437 | 0.009 | 0.630 | 0.991 | 0.663 | 23.696 |
| 0.395 | 0.010 | 0.634 | 0.990 | 0.657 | 23.484 |
| 0.335 | 0.010 | 0.637 | 0.990 | 0.648 | 23.169 |
| Model (Input) | Threshold | False Positive Rate | True Positive (Sensitivity) | Specificity | Precision | Relative Risk |
| Cancer: Pancreatic, Positive/Negative (Incidence Ratio): 1201/30382 (3.95%) | ||||||
| CATCH-FM-1b (Code Sequence) | 0.391 | 0.01 | 0.740 | 0.99 | 0.753 | 19.8 |
| XGBoost (Bag-of-words) | 0.327 | 0.01 | 0.572 | 0.99 | 0.695 | 18.3 |
| Qwen (Language Sequence) | 0.320 | 0.01 | 0.627 | 0.99 | 0.716 | 18.8 |
| Cancer: Liver, Positive/Negative (Incidence Ratio): 1099/28353 (3.88%) | ||||||
| CATCH-FM-1b (Code Sequence) | 0.508 | 0.01 | 0.696 | 0.99 | 0.736 | 19.7 |
| XGBoost (Bag-of-words) | 0.297 | 0.01 | 0.614 | 0.99 | 0.708 | 19.0 |
| Qwen (Language Sequence) | 0.239 | 0.01 | 0.591 | 0.99 | 0.696 | 18.7 |
| Cancer: Lung, Positive/Negative (Incidence Ratio): 868/30153 (2.88%) | ||||||
| CATCH-FM-1b (Code Sequence) | 0.335 | 0.01 | 0.637 | 0.99 | 0.648 | 23.2 |
| XGBoost (Bag-of-words) | 0.258 | 0.01 | 0.586 | 0.99 | 0.628 | 22.4 |
| Qwen (Language Sequence) | 0.558 | 0.01 | 0.621 | 0.99 | 0.642 | 22.9 |
Appendix F Analysis of Threshold Tuning
Clinically, models often adjust decision thresholds to prioritize sensitivity or specificity [18; 76], such as reducing false positives in prescreening or increasing sensitivity in high-risk groups. Figure 8 shows that CATCH-FM supports tunable sensitivity and specificity, which makes it well-suited for diverse clinical settings.
We evaluate our model and carefully select decision thresholds from the AUROC curve to ensure a true positive rate (TPR) 0.6, a false positive rate (FPR) 0.01, and a specificity of at least 0.99. We also report Relative Risk (RR), following suggested work in Placido et al. [65], to quantify the odds of cancer among those classified as positive compared to a random pick based just on the population disease incidence. We provide a complete threshold selection results in Table 16.
Under a fixed specificity of 0.99 and FPR of 0.01, our model demonstrates strong performance across all three cancer prediction tasks. For pancreatic cancer, a threshold of 0.39 yields a TPR of 0.74, and a relative risk of 19.8. For liver cancer, a threshold of 0.51 yields a TPR of 0.696 and a relative risk of 19.7. For lung cancer, a threshold of 0.34 yields a TPR of 0.64 and a relative risk of 23.2. These results highlight the utility of our method across diverse cancer types. Under the fixed specificity threshold, we ensure a fair comparison with the baselines, as detailed in Table 17. Across all metrics, our model consistently outperforms the baselines.
To simulate clinical implementation, following the approach proposed in Placido et al. [65], we adopt an operational decision point that simulates cost constraints, where only the top 0.1% of patients (by predicted risk) are eligible to be advanced to a surveillance program. At this threshold, our model achieves relative risk scores of 24.6, 24.9, and 35.74 for pancreatic, liver, and lung cancers prediction, respectively. This also further indicates the clinical utility of our model. We also report AUROC, AUPRC, and Relative Risk curves across thresholds to compare our model with baseline methods (XGBoost with bag-of-words and Qwen with language sequence input), as shown in Figure 9.
| Scale | Tokens | Loss | F1 | AUROC | AUPRC | Specificity | Sensitivity |
| FLOPs = 4e18, Optimal Model Scale = 260m | |||||||
| 260m | 1.7B | 2.887 | 86.6 | 97.1 | 81.2 | 99.5 | 70.7 |
| 70m | 5.2B | 2.974 | 86.4 | 97.0 | 81.1 | 99.2 | 70.7 |
| 2b | 0.2B | 3.411 | 82.7 | 96.1 | 73.3 | 99.1 | 60.9 |
| FLOPs = 2e19, Optimal Model Scale = 410m | |||||||
| 410m | 5.6B | 2.677 | 87.3 | 96.4 | 81.2 | 99.3 | 70.8 |
| 70m | 26B | 2.862 | 86.1 | 96.2 | 79.7 | 99.1 | 70.7 |
| 2b | 1.2B | 2.885 | 82.4 | 96.2 | 74.9 | 98.5 | 67.3 |
| Negative Selection | F1 (Macro) | AUROC | AUPRC | Specificity | Sensitivity | |
| Training | Testing | |||||
| Cancer: Pancreatic | ||||||
| Controlled | Controlled | 86.6 | 97.1 | 81.2 | 99.2 | 70.7 |
| Random | Random | 86.3 | 95.9 | 78.5 | 99.4 | 67.2 |
| Controlled | Random | 85.5 | 96.9 | 79.7 | 99.0 | 70.7 |
| Random | Controlled | 87.8 | 95.6 | 80.1 | 99.6 | 67.2 |
| Cancer: Liver | ||||||
| Controlled | Controlled | 84.5 | 96.0 | 76.6 | 99.3 | 63.4 |
| Random | Random | 85.0 | 95.0 | 75.8 | 99.3 | 64.5 |
| Controlled | Random | 79.4 | 94.6 | 69.0 | 98.2 | 63.4 |
| Random | Controlled | 85.2 | 94.6 | 75.5 | 99.4 | 64.5 |
| Cancer: Lung | ||||||
| Controlled | Controlled | 82.1 | 95.8 | 71.5 | 99.6 | 54.7 |
| Random | Random | 83.5 | 96.1 | 72.5 | 99.4 | 62.4 |
| Controlled | Random | 83.4 | 95.9 | 72.8 | 99.8 | 54.7 |
| Random | Controlled | 84.0 | 95.6 | 72.7 | 99.4 | 62.4 |
Appendix G Benefits of Compute-Optimal Pretraining
We show the detailed performance comparison between compute-(non-)optimal pretraining on downstream cancer pancreatic screening, as supplementary information for Figure 2, in the Table 18. We can observe that a compute-optimal model with adequate tokens outperforms larger models with insufficient tokens and smaller models with excessive tokens.
Appendix H Analysis of Cohort Control
We evaluate CATCH-FM-160m’s performance with training and testing on controlled (matched) and random (out-of-distribution) control groups to assess screening under various distribution shifts. As shown in Table 19, CATCH-FM maintains strong performance on the random control group, demonstrating its robustness in handling out-of-distribution patients and ensuring consistent results across diverse populations.
| Time Window Exclusion | AUPRC | Specificity | Sensitivity |
| Cancer: Pancreatic | |||
| 12-month | 81.2 | 99.2 | 70.7 |
| 6-month | 81.4 | 99.2 | 71.2 |
| Cancer: Liver | |||
| 12-month | 76.6 | 99.3 | 63.4 |
| 6-month | 76.9 | 99.1 | 67.7 |
| Cancer: Lung | |||
| 12-month | 71.5 | 99.6 | 54.7 |
| 6-month | 71.6 | 99.5 | 59.7 |
Appendix I Time Window Exclusion
This experiment studies CATCH-FM’s effectiveness at different exclusion window in between input medical history and cancer diagnosis. Table 20 shows the results of CATCH-FM when finetuned and evaluated with different exclusion windows. A shortened 6-month exclusion window presents an easier task, as more short term risk factors may be observable, and CATCH-FM performs better in that setup. This confirms that CATCH-FM can be conveniently adapted to different prediction settings based on healthcare professional’s preference.
| Category | Pos./Neg. (ratio) | F1 | AUROC | AUPRC | Specificity | Sensitivity |
| Cancer: Pancreatic | ||||||
| Demographics: Gender | ||||||
| Male | 616/17684 (3.48%) | 86.3 | 97.0 | 80.1 | 99.2 | 71.2 |
| Female | 584/12570 (4.65%) | 88.5 | 97.1 | 85.3 | 99.1 | 75.9 |
| Demographics: Age | ||||||
| 20-40 | 136/11018 (1.23%) | 85.9 | 93.9 | 73.8 | 99.7 | 67.8 |
| 40-60 | 434/4975 (8.72%) | 86.7 | 94.0 | 81.7 | 98.3 | 72.1 |
| 60-80 | 540/2248 (24.02%) | 86.4 | 93.7 | 87.4 | 94.8 | 77.8 |
| 80+ | 79/246 (32.11%) | 81.2 | 87.4 | 81.5 | 91.8 | 69.4 |
| Clinical Covariates: Visit Frequency | ||||||
| 100 | 627/27489 (2.28%) | 82.4 | 95.7 | 70.5 | 99.4 | 61.1 |
| 100-200 | 366/2403 (15.23%) | 90.7 | 97.6 | 92.3 | 97.4 | 84.8 |
| 200-300 | 132/374 (35.29%) | 92.7 | 98.6 | 96.8 | 95.2 | 91.4 |
| 300+ | 76/116 (65.52%) | 92.0 | 98.3 | 97.5 | 93.3 | 90.8 |
| Cancer: Liver | ||||||
| Demographics: Gender | ||||||
| Male | 690/16631 (4.15%) | 85.0 | 94.9 | 76.5 | 99.1 | 67.2 |
| Female | 408/11622 (3.51%) | 85.2 | 95.5 | 78.1 | 99.0 | 70.9 |
| Demographics: Age | ||||||
| 20-40 | 163/10072 (1.62%) | 73.4 | 87.9 | 48.2 | 99.5 | 41.3 |
| 40-60 | 441/4728 (9.33%) | 84.5 | 92.8 | 78.7 | 97.6 | 70.0 |
| 60-80 | 450/2072 (21.72%) | 88.2 | 94.2 | 88.1 | 96.1 | 79.4 |
| 80+ | 30/223 (13.45%) | 73.6 | 77.4 | 58.7 | 93.2 | 55.3 |
| Clinical Covariates: Visit Frequency | ||||||
| 100 | 566/25828 (2.19%) | 78.6 | 91.6 | 60.6 | 99.4 | 52.0 |
| 100-200 | 316/2070 (15.27%) | 89.1 | 97.7 | 90.4 | 96.7 | 83.3 |
| 200-300 | 134/360 (37.22%) | 90.0 | 97.8 | 95.0 | 91.8 | 91.8 |
| 300+ | 83/95 (87.37%) | 88.4 | 94.8 | 94.7 | 89.5 | 87.2 |
| Cancer: Lung | ||||||
| Demographics: Gender | ||||||
| Male | 498/16457 (3.03%) | 82.4 | 92.9 | 68.7 | 99.2 | 61.4 |
| Female | 369/13622 (2.71%) | 82.6 | 92.8 | 69.9 | 99.5 | 60.0 |
| Demographics: Age | ||||||
| 20-40 | 80/11654 (0.69%) | 72.7 | 78.4 | 36.2 | 99.9 | 32.3 |
| 40-60 | 293/5680 (6.16%) | 80.0 | 89.3 | 65.6 | 98.9 | 54.1 |
| 60-80 | 420/2324 (18.07%) | 82.9 | 90.5 | 78.6 | 94.8 | 70.8 |
| 80+ | 68/231 (29.44%) | 79.4 | 86.1 | 77.2 | 92.2 | 65.0 |
| Clinical Covariates: Visit Frequency | ||||||
| 100 | 401/27458 (1.46%) | 74.0 | 87.8 | 46.9 | 99.6 | 40.6 |
| 100-200 | 267/2295 (11.63%) | 86.4 | 94.8 | 83.6 | 97.2 | 75.5 |
| 200-300 | 127/314 (40.45%) | 87.8 | 95.5 | 92.4 | 91.9 | 84.7 |
| 300+ | 73/86 (84.88%) | 80.7 | 89.9 | 88.8 | 85.6 | 75.6 |
Appendix J Analysis of Demographic Subgroups
Table 21 summarizes the performance of CATCH-FM-1b across demographic subgroups. No significant bias is observed in terms of gender, as performance metrics (F1, AUROC, AUPRC) are similar between males and females across all cancers. Model performance peaks in middle-aged groups (40-60 and 60-80), with reduced sensitivity and F1 scores in younger (20-40) and older (80+) groups, likely due to these groups presenting less distinct or atypical clinical features, making detection more challenging. For the younger cohort, the limited number of positive cases may lead to data scarcity issues, resulting in reduced model sensitivity or generalizability. In contrast, the older group may exhibit more clinical complexity and comorbidities, making accurate prediction more challenging despite a higher prevalence rate. A more deep investigation is an interesting direction for future research.
| Model | FLOPs | F1 (Macro) | Accuracy | AUC | AUPRC | Specificity | Sensitivity* |
| Disease: Heart Attack, Positive/Negative (Ratio): 10205/429248 (2.38%) | |||||||
| XGBoost | - | 0.734 | 0.982 | 0.953 | 0.554 | 0.997 | 0.357 |
| LightGBM | - | 0.742 | 0.982 | 0.957 | 0.575 | 0.997 | 0.375 |
| CATCH-FM-160m | 1e18 | 0.771 | 0.981 | 0.951 | 0.579 | 0.992 | 0.508 |
| CATCH-FM-410m | 2e19 | 0.775 | 0.98 | 0.935 | 0.558 | 0.991 | 0.516 |
| CATCH-FM-720m | 4e19 | 0.777 | 0.981 | 0.935 | 0.576 | 0.992 | 0.522 |
| CATCH-FM-1b | 2e20 | 0.775 | 0.98 | 0.95 | 0.58 | 0.991 | 0.538 |
| Disease: Stroke, Positive/Negative (Ratio): 21815/625873 (3.49%) | |||||||
| XGBoost | - | 0.84 | 0.982 | 0.975 | 0.765 | 0.995 | 0.599 |
| LightGBM | - | 0.854 | 0.983 | 0.977 | 0.788 | 0.995 | 0.646 |
| CATCH-FM-160m | 1e18 | 0.873 | 0.984 | 0.982 | 0.827 | 0.994 | 0.714 |
| CATCH-FM-410m | 2e19 | 0.873 | 0.984 | 0.982 | 0.827 | 0.994 | 0.714 |
| CATCH-FM-720m | 4e19 | 0.878 | 0.985 | 0.983 | 0.834 | 0.994 | 0.727 |
| CATCH-FM-1b | 2e20 | 0.874 | 0.984 | 0.974 | 0.785 | 0.992 | 0.755 |
Appendix K Performance of Heart Attack and Stroke Screening
Building on CATCH-FM’s success in cancer screening, we extend its evaluation to other life-threatening conditions—heart attack and stroke, both leading global causes of death [60] and prevalent research topics in NHIRD [38; 10; 52; 9; 88; 86]. Early screening tools for these conditions are crucial for timely interventions. We perform binary classification to predict whether a patient will experience a stroke or heart attack exactly one year after their current clinical visit, using the preceding five years of medical history. To ensure predictions emphasize early screening efficacy, the target event (1 for detected, 0 for not detected) is strictly set at the one-year mark. Following prior work, we use ICD9 codes 43301, 43311, 43321, 43331, 43381, 43391, 43401, 43411, and 43491 for stroke [38] and the first three digits 410 for heart attack [10], applying the same case-control matching criteria as in Section 6.1. Consistent with our cancer screening approach, we evaluate CATCH-FM at compute-optimal scales and compare it against tree-based baselines.
As shown in Table 22, CATCH-FM demonstrates superior performance and outperforms all baselines with high sensitivity at 99% specificity, providing reliable early warnings for both life-threatening medical conditions. Notably, the scaling behaviors on these tasks mirror the observed scaling laws in cancer screening, further underscoring CATCH-FM’s robustness and adaptability. These findings highlight its potential for deployment in diverse, life-saving clinical scenarios. They also emphasize the value of scaling larger models with expanded medical datasets to create reliable healthcare foundation models that address critical medical use cases.
Appendix L Details of Interpretability Experiments
We process positive patients’ event token sequences using the fine-tuned CATCH-FM-1b model on each cancer dataset to obtain hidden states for every token. Following [43], we then train a TopK sparse autoencoder (SAE) on the hidden states of the [EOS] token, , which serves as an aggregated representation of patient trajectories. The SAE is implemented as follows:
where the embedding vector is passed through an encoder parameterized by and . The TopK activation function regulates the number of active latent features. The encoded representation is then reconstructed via a decoder parameterized by and . The SAE is trained using the mean squared error (MSE) loss for reconstruction.
Figure 11 presents the reconstruction evaluation across different numbers of active latent features on the cancer screening benchmark. We observe that as the number of active latent features increases, both cancer screening performance and reconstruction quality improve. Interestingly, with just 16 active latent features, the SAE’s reconstructed embeddings already capture enough information to match the original performance. This suggests that the latent cancer signal extracted from the fine-tuned CATCH-FM is inherently low-dimensional.
Appendix M Analysis of Supervised Data Scale
Many healthcare systems may not have as many patients as in NHIRD, e.g., in scattered healthcare systems. This experiment evaluates CATCH-FM with different amounts of available supervised finetuning labels. Figure 11 plots CATCH-FM’s performance finetuned with different amount of labels. CATCH-FM maintains its 99% specificity with as few as 10k training labels, with only 300 positives. Its sensitivity increases with more finetuning amount and crossed 50% with only 20k total patient data across two decades, which is fewer than a typical hospital.
Appendix N Details of Ablation Study
Table 6 evaluates the performance of various variants of CATCH-FM. Since demographics are strong risk factors for cancer, we simply use them as predictors for a clinically informed baseline. The results justifies the value of incorporating medical history features. Details are in Table 23. Moreover, the results restate the recent belief in foundation and large language models that what data to pretrain upon is more crucial than architecture twists. Changing the positional embedding and time token does not matter much, while pretraining on healthcare record boosts the sensitivity from 55.7% to 70.7%.
Training on medical code representations significantly outperforms the approach of mapping ICD codes to their text descriptions and then leveraging pretrained language models. Though the latter may bring in the advantage of additional pretraining data and shared text semantics from code names, it also brings in ambiguity and reduces the information conveyed per pretraining tokens. Our large scale NHIRD pretraining data are also sufficient to pretrain compute-optimal healthcare foundation models in the billions parameter range. Moreover, Figure 5 illustrates that pretraining in the text space tends to memorize textual phrases from medical codes rather than learning to predict the next medical events. This leads to many trivial learning tasks and ultimately results in zero training loss. How to leverage the benefits of text-based foundation models together with the conciseness of medical code modality is an interesting future research direction.
| Model | AUROC | AUPRC |
| Cancer: Pancreatic, Positive/Negative (Incidence Ratio): 1201/30382 (3.95%) | ||
| CATCH-FM-1b | 0.962 | 0.814 |
| CATCH-FM-1b (age-only) | 0.500 | 0.519 |
| XGBoost | 0.959 | 0.696 |
| XGBoost (age-only) | 0.500 | 0.519 |
| Cancer: Liver, Positive/Negative (Incidence Ratio): 1099/28353 (3.88%) | ||
| CATCH-FM-1b | 0.956 | 0.773 |
| CATCH-FM-1b (age-only) | 0.721 | 0.126 |
| XGBoost | 0.957 | 0.715 |
| XGBoost (age-only) | 0.500 | 0.519 |
| Cancer: Lung, Positive/Negative (Incidence Ratio): 868/30153 (2.88%) | ||
| CATCH-FM-1b | 0.938 | 0.691 |
| CATCH-FM-1b (age-only) | 0.712 | 0.103 |
| XGBoost | 0.953 | 0.651 |
| XGBoost (age-only) | 0.500 | 0.514 |